Find the value of n for which (8(^n+1))(2(^2n-1))(4(^n+1))/16^n=32.
Answer by scott8148(6628)

converting to a base of 2 __ (2^3(^n+1))(2(^2n-1))(2^2(^n+1))/((2^4)^n)=2^5

3n+3+2n-1+2n+2-4n=5 __ 3n+4=5 __ 3n=1